OpenStudy (anonymous):

arccos(sin(pi/6)) Help?

OpenStudy (jdoe0001):

https://www.desmos.com/calculator

OpenStudy (jdoe0001):

hmm ohhh of ... sine... ok

OpenStudy (anonymous):

@jdoe0001 I think I'm supposed to try without a calculator?

OpenStudy (jdoe0001):

http://math.info/image/325/trig_cofunction.gif <--- notice the sine cofunction so we can say that we need a value that subtracted from 1/2 gives us 1/6 that is \(\bf \cfrac{1}{2}-\theta=\cfrac{1}{6}\implies \cfrac{1}{2}-\cfrac{1}{6}=\theta\implies \cfrac{3-1}{6}=\theta\implies \cfrac{1}{3}=\theta\qquad thus\\ \quad \\ \quad \\ sin\left(\frac{\pi}{6}\right)\implies sin\left(\frac{\pi}{2}-{\color{blue}{ \frac{\pi}{3}}}\right)\implies cos\left({\color{blue}{ \frac{\pi}{3}}}\right)\qquad thus\\ \quad \\ cos^{-1}\left[sin\left(\frac{\pi}{6}\right)\right]\implies cos^{-1}\left[sin\left(\frac{\pi}{2}-{\color{blue}{ \frac{\pi}{3}}}\right)\right]\\ \quad \\ \implies cos^{-1}\left[cos\left({\color{blue}{ \frac{\pi}{3}}}\right)\right]\)

OpenStudy (jdoe0001):

see what it is?

OpenStudy (anonymous):

Not.. At.. All..

OpenStudy (jdoe0001):

well.. did you see the cofunction identities?

OpenStudy (anonymous):

Where did the 1/2 come from? And why theta?

OpenStudy (anonymous):

Yes..

OpenStudy (jdoe0001):

well, I use \(\theta\) I could have used any variable really just needed a value that when subtracted from 1/2, would give 1/6

OpenStudy (jdoe0001):

notice the sine cofunction there \(\Large \bf sin\left(\frac{\pi}{2}-{\color{blue}{ \square }}\right)=cos({\color{blue}{ \square }})\)

OpenStudy (jdoe0001):

any part you find confusing?

OpenStudy (anonymous):

All of it. So first I find 1/3. Then I.. Then it can all be equal to cos(1/3 because of those identities you showed... So then I do cos^-1(cos(pi/3))?

OpenStudy (anonymous):

BRB

OpenStudy (jdoe0001):

well... see, lemme use the angles themselves \( \bf \cfrac{\pi}{2}-\cfrac{\pi}{3}=\cfrac{3\pi-2\pi}{6}\implies \cfrac{\pi}{6}\qquad \implies sin\left(\frac{\pi}{2}-{\color{blue}{ \frac{\pi}{3}}}\right) \iff sin\left(\frac{\pi}{6}\right)\) anything confusing there?

OpenStudy (anonymous):

Okay, got you so far.

OpenStudy (anonymous):

The question asks this: Find the exact value of arccos(sin(pi/6)). For full credit, explain your reasoning.

OpenStudy (jdoe0001):

right

OpenStudy (jdoe0001):

so you see why we bother doing the \(\bf \cfrac{1}{2}-\theta=\cfrac{1}{6}\implies \cfrac{1}{2}-\cfrac{1}{6}=\theta\implies \cfrac{3-1}{6}=\theta\implies \cfrac{1}{3}=\theta\) part?

OpenStudy (anonymous):

Yes. Got it. But what do you do after you get the cos(pi/3)?

OpenStudy (jdoe0001):

say what's the \(\bf cos(\pi)\quad ?\)

OpenStudy (anonymous):

Huh?

OpenStudy (jdoe0001):

heheh, just a quick example :)

OpenStudy (anonymous):

I didn't get it.. :c

OpenStudy (jdoe0001):

as an example.... just as example... if you check your Unit Circle, what would you get for the \(\bf cos(\pi)\) ?

OpenStudy (anonymous):

Not very good at the unit circle.. 180?

OpenStudy (jdoe0001):

as an example.... just as example... if you check your Unit Circle, what would you get for the \(\bf cos(180^o)\) ?

OpenStudy (anonymous):

-1?

OpenStudy (jdoe0001):

right.... so what if take arccos() of that, what would you get from \(\bf cos^{-1}({\color{blue}{ -1}})\quad ?\)

OpenStudy (anonymous):

Oh, whoa. It's pi... cool.

OpenStudy (anonymous):

So how do I connect that to what we did?

OpenStudy (jdoe0001):

pi = 180 degrees well, what does that mean? it means \(\bf {\color{blue}{ cos(180^o)}}=-1\qquad cos^{{\color{blue}{ -1}}}(-1)=180^o\\ \quad \\\qquad \implies cos^{-1}[{\color{blue}{ cos(180^o)}}]=180^o\\ \quad \\ \quad \\ thus\implies cos^{-1}[cos(\square )]=\square \)

OpenStudy (jdoe0001):

\(\bf sin\left(\frac{\pi}{6}\right)\implies sin\left(\frac{\pi}{2}-{\color{blue}{ \frac{\pi}{3}}}\right)\implies cos\left({\color{blue}{ \frac{\pi}{3}}}\right)\qquad thus\\ \quad \\ cos^{-1}\left[sin\left(\frac{\pi}{6}\right)\right]\implies cos^{-1}\left[sin\left(\frac{\pi}{2}-{\color{blue}{ \frac{\pi}{3}}}\right)\right]\\ \quad \\ \implies cos^{-1}\left[cos\left({\color{blue}{ \frac{\pi}{3}}}\right)\right]\implies \huge ?\)

OpenStudy (anonymous):

So, afer we get 1/3, then it becomes cos^-1(cos(pi/3)). And because of the unit circle, the answer is pi/3?

OpenStudy (jdoe0001):

yeap

OpenStudy (jdoe0001):

\(\large \bf cos^{-1}\left[cos\left({\color{blue}{ \frac{\pi}{3}}}\right)\right]\implies {\color{blue}{ \frac{\pi}{3}}}\)

OpenStudy (anonymous):

I think I understand now. Thanks! :)

OpenStudy (jdoe0001):

yw